# An empirical study of validating synthetic data for formula generation

### v3

@@ Abstract
Large language models (LLMs) can be leveraged to help with writing formulas in spreadsheets, but resources on these formulas are scarce, impacting both the base performance of pre-trained models and limiting the ability to fine-tune them. Given a corpus of formulas, we can use a(nother) model to generate synthetic natural language utterances for fine-tuning. However, it is important to validate whether the natural langauge (NL) generated by the LLM is indeed accurate to be beneficial for fine-tuning. In this paper, we provide empirical results on the impact of validating these synthetic training examples with surrogate objectives that evaluate the accuracy of the synthetic annotations. We demonstrate that validation improves performance over raw data across four models (2 open and 2 closed weight). Interestingly, we show that although validation tends to prune more challenging examples, it increases the complexity of problems that models can solve after being fine-tuned on validated data.
@@
Usneek Singh Aditya Kanade Anirudh Khatry Microsoft Bangalore, India                        José Cambronero Sumit Gulwani Vu Le Mukul Singh Microsoft Redmond, US                        Gust Verbruggen Microsoft Keerbergen, Belgium
@@ 1 Introduction
Derived-column formulas in spreadsheets generate a new column by transforming existing columns in a table, and they have been shown challenging to write Gulwani et al. (2012). To aid users in writing such formulas, we can ask for a description in natural language Zhao et al. (2024). Unfortunately, since such formulas are sparse, therefore pre-trained language models (especially smaller) struggle in generating them without fine-tuning (for example, one of our models, Phi-2, achieved a pass@10 score of only 0.03, indicating a very low success rate in generating the correct formulas within 10 attempts.).
To construct a dataset for fine-tuning, public spreadsheet workbooks can be used but they contain only tables and formulas, whereas a fine-tuning dataset also requires paired natural language (NL) descriptions corresponding to each (Table, Formula). Traditionally datasets for NL-to-code tasks have been manually annotated Zhou et al. (2024); Austin et al. (2021). This is a time-consuming and expensive process. Leveraging LLMs, known for their text generation capabilities, is a viable alternative Tan et al. (2024) given that the synthetic NL generated by LLMs is accurate, as recent studies have shown that quality is more important than quantity Zhou et al. (2024); Li et al. (2023); lozhkov2024fineweb-edu.
In this paper, we leverage LLMs to predict the accuracy of synthetic NL using 3 surrogate objectives, and show empirical results of fine-tuning models on subsets of synthetic data that are accepted by these objectives. Fine-tuning models on validated subsets shows better performance in predicting formulas compared to using raw data. For example, GPT-4 fine-tuned on data validated by alternative code generation objective saw up to a 28% improvement in evaluation scores along with a 23% reduction in training time. Additionally, we observe that the models fine-tuned on validated data perform better on more complex problems. We also find that models fine-tuned on validated data still manage to learn a few functions removed during validation. Further, we release the synthetic dataset to seed future research in this area.
@@ 2 Related work
FlashFill Gulwani (2011); Gulwani et al. (2012) generates derived-column formulas by example, as users struggle with this task. SpreadsheetCoder Chen et al. (2021b) suggests formulas from surrounding context in spreadsheets. flame Joshi et al. (2024) is a small language model that understands formulas for tasks like repair and retrieval, but does not handle natural language. The NL-to-Formula (NL2F) task is introduced with a dataset obtained by converting the Text2SQL dataset to spreadsheet formulas Zhao et al. (2024). Unlike Zhao et al. (2024), our work centres on empirically evaluating different NL validation strategies.
@@
Tan et al. (2024) discusses the applications of LLMs in data annotation for classification tasks. Goel et al. (2023) demonstrates the use of LLMs in the medical domain, where they assist in labeling data with expert verification. Wang et al. (2024), Kim et al. (2024), and tang2024pdfchatannotator explore human-LLM collaborative approaches for annotation and verification. There has been no comparison of NL validation techniques on synthetic NL for NL2F.
@@ 3 Validating synthetic data
Let $T=[C_{i}]_{1}^{n}$ be a table with $n$ columns uniquely identified by $h_{i}$. A derived-column formula $F$ is a formula where each leaf node is either a constant value or a column identifier $h_{i}$. Let $U$ be an utterance in natural language that describes how to derive a column from $T$. An derived-column task is specified by $(U,T,F)$. Given $U$ and $T$ the goal is to find a formula $F^{\prime}$ such that $F^{\prime}(T)\equiv F(T)$.
@@
This validator asks the LLM to directly predict the output values $F(T)$ from $(\hat{U},T)$ and uses an element-wise row comparison to evaluate correctness. This approach leverages natural language to emulate the computation directly. It is inspired from the alternate task of output prediction discussed in Khatry et al. (2023)
@@
This validator asks the LLM to predict a program $P$ in another language (we use Python) from $(\hat{U},T)$ and compares $P(T)$ (execution of P on $T$) with $F(T)$ using element-wise comparison with the same relaxations for strings and numbers. This leverages the abilities of LLMs to generate popular programming languages Ni et al. (2023).
@@
More details about the validators are provided in Appendix 8.3.
@@ 4 Experimental setup
We mine $(T,F)$ pairs that satisfy our derived-column definition from publicly available Excel workbooks Singh et al. (2023). We create a training set and validation set of size 7833 and 422 respectively. Each $(T,F)$ pair is annotated with an utterance $\hat{U}$ using GPT-4.
@@ 5.4 RQ4: Recovery of functions removed during validation
By analyzing the output generations of the fine-tuned models, we identify new functions that were not present in their base model predictions. Our results show that the models successfully learned some of the functions that had been removed during validation (see Table 4). This suggests that fine-tuning on a high-quality dataset allows the model to remember knowledge that it had learned during pre-training, without teaching it to hallucinate on potential mistakes in the synthetic data.

### v4

@@ Abstract
Large language models (LLMs) can be leveraged to help write formulas in spreadsheets, but formula data resources are scarce, impacting both the base performance of pre-trained models and limiting the ability to fine-tune them. Given a corpus of formulas, we can use another model to generate synthetic natural language utterances for fine-tuning. However, it is important to validate whether the natural language (NL) generated by the LLM is accurate for it to be beneficial for fine-tuning. In this paper, we provide empirical results on the impact of validating these synthetic training examples with surrogate objectives that evaluate the accuracy of the synthetic annotations. We demonstrate that validation improves performance over raw data across four models (2 open and 2 closed weight). Interestingly, we show that although validation tends to prune more challenging examples, it increases the complexity of problems that models can solve after being fine-tuned on validated data.
@@
Usneek Singh Microsoft Bangalore, India                        José Cambronerothanks: Work done at Microsoft Google Atlanta, USA                        Sumit Gulwani Microsoft Redmond, USA
Aditya Kanade Microsoft Bangalore, India                        Anirudh Khatry11footnotemark: 1 University of Texas at Austin Austin, USA                        Vu Le Microsoft Redmond, USA
Mukul Singh Microsoft Redmond, USA                        Gust Verbruggen Microsoft Keerbergen, Belgium
@@ 1 Introduction
Derived-column formulas in spreadsheets generate a new column by transforming existing columns in a table, and they have been shown to be challenging to write Gulwani et al. (2012). To aid users in writing such formulas, we can ask for a description in natural language Zhao et al. (2024). Unfortunately, since such formulas are sparse, pre-trained language models (especially smaller) struggle in generating them without fine-tuning (for example, one of our models, Phi-2, achieved a pass@10 score of only 0.03, indicating a very low success rate in generating the correct formulas within 10 attempts.).
To construct a dataset for fine-tuning, public spreadsheet workbooks can be used but they contain only tables and formulas, whereas a fine-tuning dataset also requires paired natural language (NL) descriptions corresponding to each (Table, Formula). Traditionally datasets for NL-to-code tasks have been manually annotated Zhou et al. (2024); Austin et al. (2021). This is a time-consuming and expensive process. Leveraging LLMs, known for their text generation capabilities, is a viable alternative Tan et al. (2024) assuming that the synthetic NL generated by LLMs is accurate, as recent studies have shown that quality is more important than quantity Zhou et al. (2024); Li et al. (2023); Lozhkov et al. (2024).
In this paper, we leverage LLMs to predict the accuracy of synthetic NL using 3 surrogate objectives, and show empirical results of fine-tuning models on subsets of synthetic data that are accepted by these objectives. Fine-tuning models on validated subsets shows better performance in predicting formulas compared to using raw data. For example, GPT-4 fine-tuned on data validated by generating code in an alternate common programming language saw up to a 28% improvement in evaluation scores along with a 23% reduction in training time. Additionally, we observe that the models fine-tuned on validated data perform better on more complex problems. We also find that models fine-tuned on validated data still manage to learn to use functions removed during validation.
@@ 2 Related work
FlashFill Gulwani (2011); Gulwani et al. (2012) generates derived-column formulas by example, as users struggle with this task. SpreadsheetCoder Chen et al. (2021b) suggests formulas from surrounding context in spreadsheets. flame Joshi et al. (2024) is a small language model that understands formulas for tasks like repair and retrieval, but does not handle natural language. The NL-to-Formula (NL2F) task is introduced with a dataset obtained by converting the Text2SQL dataset to spreadsheet formulas Zhao et al. (2024). Unlike Zhao et al. (2024), our work centers on empirically evaluating different NL validation strategies.
@@
Tan et al. (2024) discusses the applications of LLMs in data annotation for classification tasks. Goel et al. (2023) demonstrates the use of LLMs in the medical domain, where they assist in labeling data with expert verification. Wang et al. (2024), Kim et al. (2024), and Tang et al. (2024) explore human-LLM collaborative approaches for annotation and verification. There has been no comparison of NL validation techniques on synthetic NL for NL2F.
@@ 3 Validating synthetic data
Let $T=[C_{i}]_{1}^{n}$ be a table with $n$ columns uniquely identified by a corresponding $h_{i}$ label. A derived-column formula $F$ is a formula where each leaf node in the AST (Abstract Syntax Tree) of $F$ is either a constant value or a column identifier $h_{i}$. Let $U$ be an utterance in natural language that describes how to derive a column from $T$. A derived-column task is specified by $(U,T,F)$. Given $U$ and $T$ the goal is to find a formula $F^{\prime}$ such that $F^{\prime}(T)\equiv F(T)$, where equivalence indicates both formulas produce the same outputs given the same inputs.
@@
This validator asks the LLM to directly predict the output values $F(T)$ from $(\hat{U},T)$ and uses an element-wise row comparison to evaluate correctness. For numbers, we allow an absolute difference of 0.05. For strings, we use a longest common sub-sequence ratio of 0.8 as passing criterion. This approach leverages natural language to emulate the computation directly. It is inspired from the alternate task of output prediction discussed in Khatry et al. (2023)
Figure 2: Examples of cases filtered by validators implemented on top of GPT-4. The synthetic natural language descriptions in these examples are under-specified, contain incorrect intent, or convey an unclear idea.
@@
This validator asks the LLM to predict a program $P$ in another language (we use Python) from $(\hat{U},T)$ and compares $P(T)$ (execution of P on $T$) with $F(T)$ using element-wise comparison with the same relaxations for strings and numbers previously described. This leverages the abilities of LLMs to generate popular programming languages Ni et al. (2023).
@@
More details about the validators are provided in Appendix C. We also provide a few examples in Figure 2 to illustrate cases filtered by the validators from the raw dataset.
@@ 4 Experimental setup
We mine $(T,F)$ pairs that satisfy our derived-column definition from publicly available Excel workbooks Singh et al. (2023). We create a training set and validation set of size 7833 and 422 respectively. Each $(T,F)$ pair is annotated with an utterance $\hat{U}$ using GPT-4 at a low temperature.
@@ 5.2 RQ2: Effect on fine-tuning performance
Table 3: Training time and relative improvement for different models on data subsets. Models fine-tuned on $V_{P}$ and $V_{C}$ subsets require less time than on raw data while delivering better downstream performance.
Data phi-2 mistral gpt-35 gpt-4
Raw 15h44m 8h51m 4h45m 14h00m
$V_{O}$73% -71% -60% -47%
$V_{P}$48% -45% -37% -23%
$V_{C}$36% -32% -19% -19%
@@ 5.4 RQ4: Recovery of functions removed during validation
By analyzing the output generations of the fine-tuned models, we identify new functions that were not present in their base model (without fine-tuning) predictions. Our results show that there are functions that were not in the fine-tuning dataset *and* were not in the base model’s predictions, but after finetuning on validated datasets, we see these functions used in trained model predictions (see Table 5). This suggests that fine-tuning on a high-quality dataset allows the model to remember knowledge that it had learned during pre-training, without teaching it to hallucinate on potential mistakes in the synthetic data.
@@
5.5 Recommendations
From our study, we see that a single validator (Alternate Code generation) works best on 3 out of 4 models. We also see that Output Prediction validator shows lower performance in general, likely because many functions (51) were removed from the training data during validation. However, practitioners should experiment with different validation methods, starting with the subset here, as validation in general improves performance.
@@ 7 Limitations
Appendix A Training Data Characteristics
In this section, we summarise important formula properties for the training data extracted from excel workbooks (see Table 6). From the original corpus, we remove any formulas that have deprecated functions to produce a set of 10,389 (table, formula) pairs. We then remove any pairs where the formula results in a missing/empty value for all output rows or uses multiple tables. After the process of filtering, our final dataset consists of 7,833 (table, formula) pairs. This dataset has formulas which use 122 distinct built-in functions. The most popular functions match those typically employed by Excel spreadsheet users: IF, SUM, IFERROR, CONCATENATE, AND. The other properties are summarised in Table 6). The function call count refers to the frequency of Excel function calls within a formula. The depth of formulas denotes the extent of nested function calls within them. Operator count is the number of arithmetic operators (+, -, *, /) in a formula.
Table 6: Characteristics of formulas used in Training Data obtained from Excel spreadsheets
Fxn. call count Formula depth Op. count
0 3554 3554 2887
1 2625 2682 2811
2 965 1030 1169
3 285 325 435
4 115 125 187
$\geq 5$ 289 113 344
Appendix B Model hyper-parameters used while Fine-tuning
Phi-2
For the Phi-2 model, fine-tuning was performed for 10 epochs with a batch size of 8. The learning rate was set to 1e-6, and the Adam optimizer was used along with a cross-entropy loss function.
Mistral
The Mistral model was fine-tuned for 15 epochs using the LoRA technique Hu et al. (2021). The specific parameters for LoRA included a LoRA rank ($Lora\_r$) of 64, a LoRA alpha ($Lora\_alpha$) of 16, and a LoRA dropout ($Lora\_dropout$) of 0.1. The target modules for LoRA adaptation were "q_proj", "k_proj", "v_proj", "o_proj", "gate_proj", "up_proj", "down_proj", and "lm_head". No bias configuration was used, and the task type was Causal Language Modeling (CAUSAL_LM). The learning rate for this model was set to 2e-4, and the batch size was 8. Optimization was carried out using the PagedAdamW 32-bit optimizer.
GPT-35 and GPT-4
They have been fine-tuned using LoRa on default settings used for these models in Azure API documentation [^1] https://learn.microsoft.com/en-us/azure/ai-services/openai/.
Appendix C Technical Details of validators
In this section, we provide the details about the prompts used with each validator in the above study. We use greedy decoding for all prompts to ensure more precise computation.
Output Prediction
We use an LLM (GPT-4) as a validator V. We prompt the LLM with input table and NL to compute the output for the target column directly. Then we validate the target output by comparing them with the actual outputs, with validation deemed successful only if the expected and actual outputs match for all rows in the table. The matching criteria differs based on the datatype: for a numeric value we allow an absolute difference of up to 0.05 and a string is considered a match when the longest matching contiguous sub-sequence coefficient (defined as length of longest matching sub-sequence divided by length of the longer string) is greater than 0.8. The prompt used for this technique is provided in Figure 5.
Figure 5: Prompt used for Output Prediction validation
Alternate code generation
We use an LLM (GPT-4) as V and task it with Python generation using NL and table as the input. The matching criterion is same as that of Direct computation. The prompt is provided in Figure 6.
Figure 6: Prompt used for Alternate code generation validation
Classification
We prompt an LLM (GPT-4) as validator V to generate a binary outcome, judging whether the given natural language query accurately describes the formula when applied to the corresponding table. The prompt is provided in Figure 7.
Figure 7: Prompt used for Classification validation